Amy travels 450 miles at a certain speed. If the car had gone 15mph faster, the trip would have taken one hour less. Find Amy's speed.
A. 78 mph
B. 75 mph
C. 68 mph
D. 72 mph

Answer by venugopalramana(3286)

LET AMY'S SPEED BE X MPH….TIME TO TRAVEL 450 MILES=450/X HRS.
IF CAR GOES 15 MPH FASTER SPEED = X+15…..TIME TO TRAVEL=450/(X+15)
THIS TIME IS 1 HOUR LESS.SO…450/X-450/(X+15)=1…MULTIPLYING BY LCM X(X+15)
450{X+15-X)=X(X+15)=450*15=6750=X^2+15X
X^2+15X-6750=0=X^2+90X-75X-6750=X(X+90)-75(X+90)=(X-75)(X+90).HENCE X=75….B IS THE ANSWER.